Suppose you are a salesperson working 50 weeks per year, five days a week, eight hours a day. You want to make $50,000 per year, which is based on a 10 percent commission of gross sales. How many hours of face time with customers can you expect in any given year (you can make an assumption like 40% time for face to face selling and 60% for non-selling activities)? How much will you have to generate in sales per hour to make $50,000?
Answer #1

Answer: face time with customers= 40% of 50 weeks*5 days*8 hours = 0.4*50*8*5= 800 hours

Sales to make $50,000= 0.1*revenue

Revenue= 50,000/0.1= 500,000

sales per hour to make $50,000= 500,000/face time with customers = $625